# When and How Does CLIP Enable Domain and Compositional Generalization?

### v1

@@ Appendix B Additional Results For Section 5
In Figure 2, we compared the effective robustness trends of our different training setup across different domains. Table 6 shows the final test performances (averaged over three runs) of our CLIP models with ImageNet-Captions as the base dataset and ResNet-50 as the vision encoder.
Table 6: CLIP robustness results for the considered experimental conditions (Figure 1). We repeated CLIP trainings three times and evaluated the models on the unseen classes of the test domain. CLIP generalizes better with higher domain diversity. Intriguingly, CLIP achieves superior performance when not seeing the domain at all vs. seeing a subset of it. However, while diversity substantially improves CLIP’s generalization performance, there remains a performance gap to a model that has seen similar samples to the classes $C_{2}$. Figure 2 shows the respective effective robustness plots.
@@
Figure 3 shows which domains are the most helpful to include for generalizing to sketch images, and reveals that clipart and painting contribute strongly, while infograph and quickdraw have little benefit or even slightly decrease performance. We also conducted the same experiment with clipart as the test domain. Figure 8 shows that sketches are also the most helpful domain to include for generalizing to clipart, suggesting that both domains might leverage similar visual features. The inclusion order of the remaining domains remained the same as in the sketch experiment with the added difference that after including painting, both infograph and quickdraw slightly decreased generalization performance.
Figure 8: Domain-wise robustness gains for clipart. Figure 3 shows that clipart is the most helpful domain to include for generalizing to sketches. Similarly, sketches are also the most helpful for generalizing to clipart.
@@ B.1 Validity of the Results across Architecture, Dataset, and Loss Choices
We conducted the experiments from the main text (Figures 2 and 6) using ImageNet-Captions as the base dataset and a ResNet-50 vision encoder. To ensure the consistency of our findings across different base datasets, vision encoder architectures, and contrastive loss functions, we performed additional experiments by systematically varying each of these components. Due to computational resource constraints, we repeated these experiments only for the clipart and sketch domains, where increasing domain diversity yielded the most significant robustness gains.
@@
Tables 7 and 9 confirm that for transformer-based vision encoders, robustness also improves with increasing domain diversity, as expected. Similarly, compositional generalization performs worse than domain generalization.
Table 7: Results when varying CLIP’s vision encoder. We find similar trends across these vision encoder choices.
@@
For the base dataset experiments, we trained our ResNet-50 CLIP configuration using CC3M and CC12M as the base datasets. Following Radford et al. (2021), we used a maximum learning rate of 5e-4. We further adjusted the number of warmup steps to 2000 and the batch size to 2048. All other hyperparameters were consistent with the ResNet-50-based experiments (Section A.6).
Tables 8 and 10 show that both CC3M and CC12M exhibit the same robustness trends as our ImageNet-Captions models. However, the robustness gains are smaller compared to ImageNet-Captions. This is most likely due to the reduced relative weighting of the DomainNet images with larger base dataset sizes, as well as the higher inherent diversity of CC3M and CC12M. For example, both CC3M and CC12M include some non-natural images, which may diminish the effect of further increasing domain diversity. Note that compositional generalization seems to be slightly better than domain generalization for the larger datasets, which we attribute to domain contamination (see Section B.2 why this may benefit compositional generalization). We leave further investigation into the effect of scale and diversity of the base dataset for future work.
Table 8: Results when varying the base dataset $D_{0}$. We observe similar trends as for CC3M and CC12M as for ImageNet-Captions. The only exception is that compositional generalization now tends to always work better than domain generalization for CC12M. We attribute this to a domain contamination of CC12M, i.e., CC12M contains a lot of sketch and clipart images.
@@
For the loss experiments, we trained our ResNet-50 CLIP configuration but replaced the standard CLIP loss with the SigLIP loss (Zhai et al., 2023).
Tables 9 and 11 show that our SigLIP models are consistent with our observations from the experiments in the main text.
@@ B.2 Challenges of Compositional Generalization
In practice, however, enforcing little to no class overlap may not always be feasible; particularly in zero-shot settings where CLIP is applied to data and/or tasks that are unknown at training time. Therefore, we investigated the severity of this in greater detail. To do this, we partitioned the set of classes $C=C_{1}\cup C_{2}\cup C_{3}$ in our test domain $D_{i}$ into three disjoint subsets (Figure 12(a)):
@@
For this experiment, we selected the sketch domain as our test domain. [^4] We chose the sketch domain due to the availability of a large class distribution from ImageNet-Sketch. The subsets $C_{1}$ and $C_{2}$ were defined as described in Section 3, meaning that $C_{1}\cup C_{2}$ represents the class distribution queried during evaluation (i.e., all DomainNet classes). For $C_{3}$, we used classes from ImageNet-Sketch (Wang et al., 2019) that do not overlap with the classes from DomainNet (see Section A.7 for more details). Note that throughout all these experiments, the class distribution in the other domains $D_{j\neq i}$ remained unchanged–that is, classes from $C_{1}$ and $C_{2}$ were included in training, while no classes from $C_{3}$ were introduced.
As shown in Tables 2 and 12(b) (rightmost bar), training exclusively on samples from $C_{3}$ significantly improves compositional generalization performance compared to not seeing any test domain samples from $D_{i}$ (zero line) or only seeing samples from classes $C_{1}$ of $D_{i}$ (leftmost bar). To further investigate this, we trained CLIP models using samples from $C_{1}\cup C_{3}$ of the test domain $D_{i}$ and gradually reduced the number of classes from $C_{1}$, while keeping all other factors fixed. Figure 12(b) shows that reducing class overlap (moving from left to right)–i.e., decreasing the number of $C_{1}$ classes seen during training–consistently improves compositional generalization performance. This reaffirms the detrimental effect of class overlap, as demonstrated in Table 2, but also highlights that increasing class diversity can help mitigate its impact on compositional generalization performance.
@@
Figure 12: Severity of class overlap on compositional generalization. 12(a): We partitioned the classes $C$ in the test domain $D_{i}$ into three disjoint subsets: $C=C_{1}\cup C_{2}\cup C_{3}$. During evaluation, both $C_{1}$ and $C_{2}$ are included in the query set, allowing the model to predict any class from these subsets. However, the actual test set, $D_{i}^{C_{2}}$, only contains samples from $C_{2}$. In contrast, the classes in $C_{3}$ are excluded from both the query and test sets. This partitioning allows us to investigate the severity of class overlap for compositional generalization performance. In particular, we found that including classes of $C_{1}$ led to a significant drop in compositional generalization–even performing worse than domain generalization (Figure 2). On the other hand, replacing the class samples from $C_{1}$ with the ones of $C_{3}$ (which are not part of the query set) resulted in a significant improvement in compositional generalization (Table 2). 12(b): To investigate the severity, we varied the mixture of classes from $C_{1}$ and $C_{3}$. We find that reducing classes from $C_{1}$ (moving from left to right), while keeping all other factors fixed, steadily improves compositional generalization performance (with the exceptions for the class mixtures 25% $C_{1}$ + 100% $C_{3}$ and 1 class $C_{1}$ + 100% $C_{3}$).
@@ B.3 Closing The Generalization Gap
For a given test domain $D_{i}$, we considered the number of samples of our test set $D_{i}^{C_{2}}$ to be 100% and then trained additional CLIP models in which we successively added 5%, 10%, 15%, 20%, 40%, 60%, and 80% of these samples to the training data. Note that we ensured that the overall dataset size remained unchanged.
Figure 13 shows that performance on the classes seems to follow a roughly linear relationship with the number of samples allowed for all domains, except for the quickdraw domain. For quickdraw, the relationship seems to be log-linear instead, which may be due to the fact that CLIP does not generalize at all for quickdraw. This observation also relates to the findings of Udandarao et al. (2024), who predict that a linear increase in samples leads only to a log-linear increase in zero-shot performance. Across all domains and settings, we observed that to achieve the maximally possible performance, 100% of the samples from $D_{i}^{C_{2}}$ are required.
@@
Figure 13: Interpolation between CG settings with and without domain-specific samples of the test classes $C_{2}$. For each interpolation, we fitted both a linear and a log-linear regression model and visualized the fit with a lower mean squared error (MSE). Performance on the test classes appears to follow a roughly linear relationship with the number of test samples included, except for quickdraw, which shows a strong log-linear relationship.
@@ B.4 Role of Language Supervision
Tables 11 and 14 show the results of our experiments on supervised classifiers. We found that CLIP models consistently exhibit slightly higher robustness than their supervised counterparts, which may be due to the richness of captions (Xue et al., 2024; Wen et al., 2024). Interestingly, CLIP’s advantages are more pronounced in compositional generalization settings (CG low/high-diversity). However, in the domain generalization setting (Leave-out-domain), supervised classifiers can sometimes achieve generalization comparable to CLIP.
Table 11: Performance comparison across supervised experiments. The robustness of supervised models also increases with domain diversity. However, supervised classifiers generalize slightly worse than CLIP models (Figures 4 and 14).
@@ Appendix C Additional Technical Details and Results for Section 6
C.1 Technical Details on Sparse Autoencoders and Additional Results
@@ C.2 Technical Details on Center Kernel Alignment and Additional Results
We computed center kernel alignment between all pairs of domains for each class. Thereby, we can assess the representational similarity for each class across the domains. For visualization, we averaged over all classes to obtain a representational similarity estimate for each domain. In the main text, we further averaged together the non-quickdraw domains (all domains are shown in Figures 15 and 16).
@@
Figures 15 and 16 show results for all classes $C$, the classes seen during training $C_{1}$, and the unseen classes $C_{2}$. Interestingly, we find that the representational gap widens for the unseen classes $C_{2}$.
@@
Figure 15: Linear center kernel alignment (CKA) similarity. Quickdraw has the lowest representational similarity across domains. This particularly emphasized for the unseen classes $C_{2}$ (15(c), 15(f)).
@@
Figure 16: Non-linear center kernel alignment (CKA) similarity. The results follow the same pattern as the linear CKA (Figures 6(a) and 15): Quickdraw is the most representational dissimilar domain.
@@ C.3 Technical Details on the Circuit Similarity Analysis and Additional Results
Attributing the causal effects of model components is a key goal of (mechanistic) interpretability research (Meng et al., 2022; Marks et al., 2024; Mueller et al., 2024). We drew inspiration from them to analyze the level of sharing of the most important CLIP’s model components, i.e., the axis-aligned neurons in its vision encoder, across domains. To do this, we first must attribute their importance via the indirect effect (Pearl, 2001).
@@
Since there are lot of neurons in CLIP’s vision encoder, we sped up the computation through a linear approximation using integrated gradients (Sundararajan et al., 2017), following Marks et al. (2024):
@@
where $\alpha\in\{0,\frac{1}{N},\cdots,\frac{N-1}{N}\}$) and we set $\mathbf{a}^{l,n}_{\text{patch}}$ to $\mathbf{0}$. We adapted the codebase from Marks et al. (2024) (https://github.com/saprmarks/feature-circuits, License: MIT) for our implementation.
@@
We compared the circuit similarity of pairs of graphs resorting to graph similarity measures. For example, we computed the Jaccard index for the nodes $N_{C_{\text{domain}~{}d_{1}\text{, class c}}},\;N_{C_{\text{domain}~{}d_{2}\text{, class c}}}$, as follows:
@@
However, the simple node overlap cannot capture more complex structural and hierarchical similarities between circuits. Thus, we used the normalized Weisfeiler-Lehman subtree graph kernel (Shervashidze et al., 2011). The Weisfeiler-Lehman graph kernel is popular graph kernel choice since (1) it can handle labeled, directed graphs of different sizes, (2) it is expressive, and (3) scales well to large graphs. The main idea of the Weisfeiler-Lehman kernel is the following procedure, given two labeled graphs, $G_{1}$ and $G_{2}$:
@@
Figures 17 and 12 shows results for all classes $C$, the classes seen during training $C_{1}$, and the unseen classes $C_{2}$. Interestingly, we find that, similar as for the representational similarities in Figures 15 and 16, nodes are less shared and circuits are slightly less similar for the unseen classes.
@@
Figure 17: Amount of shared neurons. The quickdraw domains shares the least neurons. This is especially apparent for the unseen classes $C_{2}$ (17(c), 17(f)).
@@
Table 12: Weisfeiler-Lehman similarities. Higher similarities mean higher circuit (graph) similarity. The quickdraw domain exhibits the least degree of similarity, supporting our hypothesis that sharing of the circuitry is critical for generalization.

### v2

@@ Appendix B Additional Results For Section 5
In Figure 2, we compared the effective robustness trends of our different training setup across different domains. Table 6 shows the final balanced top-1 test accuracy (averaged over three runs) of our CLIP models with ImageNet-Captions as the base dataset and ResNet-50 as the vision encoder.
Table 6: CLIP robustness results for the four training data setups (see Figure 1). We repeated CLIP trainings three times and evaluated the model’s balanced top-1 accuracy on the unseen classes of the test domain. CLIP generalizes better with higher domain diversity. Intriguingly, CLIP achieves superior performance when not seeing the domain at all vs. seeing a subset of it (c.f., Table 2). However, while diversity substantially improves CLIP’s generalization performance, there remains a performance gap to a model that has seen similar samples to the classes $C_{2}$ (c.f., Table 1). Figure 2 shows the respective effective robustness plots.
@@
Other evaluation metrics
We also evaluated our CLIP models using balanced top-5 accuracy and macro $F_{1}$ score to ensure our findings remain consistent across different evaluation metrics. Figure 8 shows the corresponding effective robustness plots.
Figure 8: Effective robustness for additional evaluation metrics. We re-visualized the results from Figure 2 using balanced top-5 accuracy (first row) and macro $F_{1}$ score (second row). Our findings remain consistent across both metrics.
@@ B.1 Validity of the Results across Architecture, Dataset, and Loss Choices
We conducted the experiments from the main text (Figures 2 and 6) using ImageNet-Captions as the base dataset and a ResNet-50 vision encoder. To ensure the consistency of our findings across different base datasets, vision encoder architectures, and contrastive loss functions, we performed additional experiments by systematically varying each of these components. In addition, we also investigated BLIP (Li et al., 2022), a different method for vision-language pre-training. Due to computational resource constraints, we repeated these experiments only for the Clipart and Sketch domains, where increasing domain diversity yielded the most significant robustness gains.
@@
Tables 7 and 9 confirm that for transformer-based vision encoders, robustness also improves with increasing domain diversity, as expected. Similarly, compositional generalization can perform worse than domain generalization.
Table 7: Results (balanced top-1 accuracy) when varying CLIP’s vision encoder. We find similar trends across these vision encoder choices.
@@
For the base dataset experiments, we trained our ResNet-50 CLIP configuration using CC3M and CC12M as the base datasets. Following Radford et al. (2021), we used a maximum learning rate of 5e-4. We also adjusted the number of warmup steps to 2000 and the batch size to 2048. All other hyperparameters were consistent with the experiments from the main text (see Section A.6 for the hyperparameters).
Tables 8 and 10 confirm that both CC3M and CC12M exhibit the same robustness trends as our ImageNet-Captions models. However, the robustness gains are smaller compared to ImageNet-Captions. This is most likely due to the reduced relative weighting of the DomainNet images with larger base dataset sizes, as well as the higher inherent diversity of CC3M and CC12M. For example, both CC3M and CC12M include some non-natural images, which may diminish the effect of further increasing domain diversity. Note that compositional generalization seems to be slightly better than domain generalization for larger datasets, which we attribute to domain contamination (see Section B.2 why this may benefit compositional generalization). We leave further investigation into the effect of dataset size and diversity of the base dataset for future work (see Section 7 for a discussion).
Table 8: Results when varying the base dataset $D_{0}$. We observe similar trends for CC3M and CC12M as for ImageNet-Captions. The only exception is that compositional generalization now tends to always work better than domain generalization for CC12M. We attribute this to a domain contamination of CC12M, i.e., manual inspection shows that CC12M contains a lot of Sketch and Clipart images.
@@
For the loss experiments, we trained our ResNet-50 CLIP configuration but replaced the original CLIP loss with the SigLIP loss (Zhai et al., 2023).
Tables 9 and 11 confirm that the results for models trained with SigLIP are consistent with our observations from the experiments in the main text.
@@
Beyond Contrastive Losses
[…]
Table 10 and Figure 12 show that the results for BLIP follow the same trends as our CLIP models, except for the overall higher performance due to the pre-trained image and text encoders.
Table 10: Results when using BLIP (Li et al., 2022). Using BLIP as a pre-training method does not change the trends observed for standard CLIP. Note that the overall increased performance of BLIP is due to their usage of pre-trained image and text encoders.
Training regime Training data setup Clipart Sketch
CLIP Natural-only 19.4 19.6
Leave-out-domain 27.1 31.8
CG low-diversity 18.7 16.4
CG high-diversity 28.6 28.6
BLIP Natural-only 37.1 35.7
Leave-out-domain 61.8 58.6
CG low-diversity 35.0 41.6
CG high-diversity 55.6 61.2
Figure 12: Effective robustness plots for BLIP. Refer to Figure 2 for effective robustness plots when using standard CLIP.
@@ B.2 Challenges of Compositional Generalization
In practice, however, enforcing little to no class overlap may not always be feasible; particularly in zero-shot settings where CLIP is applied to data and/or tasks that are unknown at training time. Therefore, we investigated the severity of this in more detail. To do this, we partitioned the set of classes $C=C_{1}\cup C_{2}\cup C_{3}$ in our test domain $D_{i}$ into three disjoint subsets (see Figure 13(a)):
@@
For this experiment, we selected Sketch as our test domain. [^5] We chose Sketch due to the availability of a large class distribution from ImageNet-Sketch. The subsets $C_{1}$ and $C_{2}$ were defined as described in Section 3, meaning that $C_{1}\cup C_{2}$ represents the class distribution queried during evaluation (i.e., all DomainNet classes). For $C_{3}$, we used classes from ImageNet-Sketch (Wang et al., 2019) that do not overlap with the classes from DomainNet (see Section A.3 for more details). Note that throughout all these experiments, the class distribution in the other domains $D_{j\neq i}$ remained unchanged—that is, classes from $C_{1}$ and $C_{2}$ were included in training, while no classes from $C_{3}$ were introduced.
As shown in Tables 2 and 13(b) (rightmost bar), training solely on samples from $C_{3}$ significantly improves compositional generalization performance compared to not seeing any test domain samples from $D_{i}$ (zero line) or only seeing samples from classes $C_{1}$ of $D_{i}$ (leftmost bar). To further investigate this, we trained CLIP models using samples from $C_{1}\cup C_{3}$ of the test domain $D_{i}$ and gradually reduced the number of classes from $C_{1}$, while keeping all other factors fixed. Figure 13(b) shows that reducing class overlap (moving from left to right)—i.e., decreasing the number of $C_{1}$ classes seen during training—consistently improves compositional generalization performance. This reaffirms the detrimental effect of class overlap, as observed in Table 2, but also highlights that increasing class diversity can help mitigate its impact on compositional generalization performance.
@@
Figure 13: Severity of class overlap on compositional generalization. 13(a): We partitioned the classes $C$ in the test domain $D_{i}$ into three disjoint subsets: $C=C_{1}\cup C_{2}\cup C_{3}$. During evaluation, both $C_{1}$ and $C_{2}$ are included in the query set, allowing the model to predict any class from these subsets. However, the actual test set, $D_{i}^{C_{2}}$, only contains samples from $C_{2}$. In contrast, the classes in $C_{3}$ are excluded from both the query and test sets. This partitioning allows us to investigate the severity of class overlap for compositional generalization performance. In particular, we found that including classes of $C_{1}$ led to a significant drop in balanced top-1 accuracy for compositional generalization—even performing worse than domain generalization (Figure 2). On the other hand, replacing the class samples from $C_{1}$ with the ones of $C_{3}$ (which are not part of the query set) resulted in a significant improvement in compositional generalization (Table 2). 13(b): To investigate the severity, we varied the mixture of classes from $C_{1}$ and $C_{3}$. We find that reducing classes from $C_{1}$ (moving from left to right), while keeping all other factors fixed, steadily improves compositional generalization performance (with the exceptions for the class mixtures 25% $C_{1}$ + 100% $C_{3}$ and 1 class $C_{1}$ + 100% $C_{3}$).
@@ B.3 Closing The Generalization Gap
For a given test domain $D_{i}$, we consider the number of training samples of our test set $D_{i}^{C_{2}}$ to be 100% and then trained additional CLIP models in which we successively added 5%, 10%, 15%, 20%, 40%, 60%, and 80% of training samples from $D_{i}^{C_{2}}$ to the training data. Note that we ensured that the overall dataset size remained unchanged.
Figure 14 shows that performance on the classes seems to follow a roughly linear relationship with the number of samples allowed for all domains, except for the Quickdraw domain. For Quickdraw, the relationship seems to be log-linear instead, which may be due to the fact that CLIP does not generalize at all for Quickdraw. This observation also relates to the findings of Udandarao et al. (2024), who predict that a linear increase in samples leads only to a log-linear increase in zero-shot performance. Across all domains and settings, we observed that to achieve the maximally possible performance, 100% of the training samples from $D_{i}^{C_{2}}$ are required.
@@
Figure 14: Interpolation between CG settings with and without domain-specific training samples of the test classes $C_{2}$. For each interpolation, we fitted both a linear and a log-linear regression model and visualized the fit with a lower mean squared error (MSE). Balanced top-1 accuracy on the test classes appears to follow a roughly linear relationship with the number of test samples included, except for Quickdraw, which shows a strong log-linear relationship.
@@ B.4 Role of Language Supervision
Tables 12 and 15 show the results of our experiments on supervised classifiers. We found that CLIP models consistently exhibit slightly higher robustness than their supervised counterparts, which may be due to the richness of captions (Xue et al., 2024; Wen et al., 2024). Interestingly, CLIP’s advantages are more pronounced in compositional generalization settings (CG low/high-diversity). However, in the domain generalization setting (Leave-out-domain), supervised classifiers can sometimes achieve generalization comparable to CLIP.
Table 12: Performance comparison across supervised experiments. The robustness of supervised models (as measured by balanced top-1 accuracy) also increases with domain diversity. However, supervised classifiers generalize slightly worse than CLIP models (Figures 4 and 15).
@@ Appendix C Additional Technical Details and Results for Section 6
C.1 Technical Details on Sparse Autoencoders
@@
We trained one SAE for our natural-only baseline, two SAE’s for the leave-out-domain setting (with Clipart and Sketch as the test domain), and another two SAEs for the CG high-diversity setting (also with Clipart and Sketch as test domain). For each model, we analyzed the extent to which the top-k most activating SAE features are shared between the test domain (Clipart or Sketch) and all other domains (not just the training domains).
To identify the top-k SAE features, we computed the SAE hidden representations (after applying the non-linearity $\phi$) for each sample of each class and domain. We then selected the top-k SAE features (with $k\in\{5,10,15,20\}$) per class-domain pair based on how frequently they ranked among the top-20 most activating features, i.e., largest activation magnitudes (see Algorithm 1).
To measure feature sharing, we then calculated the percentage overlap between the top-k SAE features of the test domain and those of each of the other domains. To yield a single overlap score per model, we averaged across classes $C_{2}$, these pairs of domains, and the four values of $k$ (see Algorithm 2). Finally, we calculated the increase of the overlap score of the corresponding leave-out-domain and CG high-diversity models over the natural-only baseline model.
Algorithm 1 get_topk_features
0:  domain of interest $D_{i}$, class of interest $c$, number of features to consider $k$
1:  $\text{hist}\leftarrow\mathbf{0}\in\mathbb{R}^{h}$ // $h$ is the number of concepts
2:  for each image $I\in D_{i}^{c}$ do
3:     $\mathrm{a}\leftarrow f_{\text{CLIP}}(I)$ // CLIP image activation
4:     $w\leftarrow\phi(f_{\text{SAE}}(\mathrm{a}))$ // SAE concept activations
5:     for each index $i\in\text{topk}(w,20)$ do
6:        $\text{hist}[i]\leftarrow\text{hist}[i]+1$ // count top-20 activating concepts
7:     end for
8:  end for
8:  topk(hist, $k$) // return top-k most frequent features
Algorithm 2 measure_feature_sharing
0:  test domain $D_{i}$
1:  $S\leftarrow\emptyset$ // initialize empty set of scores
2:  for each $k\in\{5,10,15,20\}$ do
3:     for each class $c\in C_{2}$ do
4:        $F_{i}\leftarrow\text{get\_topk\_features}(D_{i},c,k)$ // top-k features of test domain $D_{i}$
5:        for each other domain $D_{j}\neq D_{i}$ do
6:           $F_{j}\leftarrow\text{get\_topk\_features}(D_{j},c,k)$ // top-k features of domain $D_{j}$
7:           $s\leftarrow\frac{|F_{i}\cap F_{j}|}{k}$ // percentage overlap of $F_{i}$ and $F_{j}$
8:           $S\leftarrow S\cup\{s\}$
9:        end for
10:     end for
11:  end for
11:  $\frac{1}{|S|}\sum\limits_{s\in S}s$ // overlap averaged over domains $D_{j}$, classes $c$, and $k$
@@ C.2 Technical Details on Center Kernel Alignment and Additional Results
We computed center kernel alignment between all pairs of domains for each class. Thereby, we can assess the representational similarity for each class across the domains. For visualization, for each domain, we averaged over all classes to obtain a representational similarity estimate. In the main text, we further averaged together the non-Quickdraw domains (all domains are shown in Figures 16 and 17).
@@
Figures 16 and 17 show results for all classes $C$, the classes seen during training $C_{1}$, and the unseen classes $C_{2}$. Interestingly, we find that the representational gap widens for the unseen classes $C_{2}$.
@@
Figure 16: Linear center kernel alignment (CKA) similarity. Quickdraw has the lowest representational similarity across domains. This particularly emphasized for the unseen classes $C_{2}$ (16(c), 16(f)).
@@
Figure 17: Non-linear center kernel alignment (CKA) similarity. The results follow the same pattern as the linear CKA (Figures 6(a) and 16): Quickdraw is the most representational dissimilar domain.
@@ C.3 Technical Details on the Circuit Similarity Analysis and Additional Results
Attributing the causal effects of model components is a key goal of (mechanistic) interpretability research (Meng et al., 2022; Marks et al., 2025; Mueller et al., 2024). We drew inspiration from them to analyze the level of sharing of the most important CLIP’s model components, i.e., the axis-aligned neurons in its vision encoder, across domains. To do this, we first must attribute the components’ importance via the indirect effect (Pearl, 2001).
@@
Since there are lot of neurons in CLIP’s vision encoder, we sped up the computation through a linear approximation using integrated gradients (Sundararajan et al., 2017), following Marks et al. (2025); Hanna et al. (2024):
@@
where $\alpha\in\{0,\frac{1}{N},\cdots,\frac{N-1}{N}\}$) and we set $\mathbf{a}^{l,n}_{\text{patch}}$ to $\mathbf{0}$. We adapted the codebase from Marks et al. (2025) (https://github.com/saprmarks/feature-circuits, License: MIT)—using nnsight (Fiotto-Kaufman et al., 2025, https://github.com/ndif-team/nnsight, License: MIT) internally—for our implementation.
@@
We compared the circuit similarity of pairs of graphs resorting to classical graph similarity measures.
Node-level similarity
We computed the Jaccard index for the nodes $N_{C_{\text{domain}~{}D_{1}\text{, class c}}},\;N_{C_{\text{domain}~{}D_{2}\text{, class c}}}$, as follows:
@@
Intuitively, this quantifies the degree of overlap between the most contributing model components for class $c$ across domains $D_{1}$ and $D_{2}$.
Graph-level (structural) similarity
Above simple node overlap cannot capture more complex structural and hierarchical similarities between circuits. Thus, we used the normalized Weisfeiler-Lehman subtree graph kernel (Shervashidze et al., 2011). The Weisfeiler-Lehman graph kernel is popular graph kernel choice since (1) it can handle labeled, directed graphs of different sizes, (2) it is expressive, and (3) scales well to large graphs. The main idea of the Weisfeiler-Lehman kernel is the following procedure, given two labeled graphs, $G_{1}$ and $G_{2}$:
@@
False positives
In certain cases, the Weisfeiler-Lehman kernel can incorrectly identify non-isomorphic graphs as isomorphic. However, several factors reduce the chance of such false positives in our setting. Specifically, we used $h=3$, which yields a more expressive graph kernel. In addition, node names in the circuits encode global topological information (i.e., layer and neuron indices), which further reduces the chance of false positives. Finally, we verified that the circuits were already distinguishable at the node level (i.e., at the 0-th iteration of the Weisfeiler-Lehman kernel). [^7] False positives can only occur if two graphs remain indistinguishable across all iterations of the Weisfeiler-Lehman kernel.
@@
Figures 18 and 13 shows results for all classes $C$, the classes seen during training $C_{1}$, and the unseen classes $C_{2}$. Interestingly, we find that, similar as for the representational similarities in Figures 16 and 17, nodes are less shared and circuits are slightly less similar for the unseen classes $C_{2}$.
@@
Figure 18: Amount of shared neurons. The Quickdraw domains shares the least neurons. This is especially apparent for the unseen classes $C_{2}$ (18(c), 18(f)).
Table 13: Weisfeiler-Lehman similarities. Higher similarities mean higher circuit (graph) similarity. The Quickdraw domain exhibits the least degree of similarity, supporting our hypothesis that sharing of the circuits is critical for generalization.